# Dual Test-time Training for Out-of-distribution Recommender System

### v1

@@ V-A Datasets & Metric
V-B Experiment Settings
V-B1 Implementation Details.
The experiments are executed utilizing an NVIDIA 3090 GPU on the PyTorch platform. For the benchmark models, the performance values from the COR [11] paper are directly reported for all datasets. In our proposed approach, we utilized COR as the primary network and trained our network using the Adam optimizer [28]. The batch size was established at 500. Moreover, the learning rate and dropout rate were set at 1e-4 and 0.5 respectively for all datasets. The training during test-time was carried out for 10 epochs. The trade-off hyper-parameter $\alpha$ was selected from the set $\{0.01,0.1,1.0,10,100\}$. The number of clusters $K$ was consistently set at 4, and the threshold for selecting high confidence samples was maintained at 0.9.
V-B2 Comparison Recommendation Methods.
To substantiate the efficiency of our DT3OR, we contrast its performance with several baseline models, including FM [29], NFM [30], MultiVAE [31], MacridVAE [32], MacridVAE+FM, CausPref [12], and COR [11]. Specifically,
@@ V-B Experiment Settings
V-C (RQ1) Performance Comparisons
We evaluate the experiments against seven baseline models based on Recall and NDCG metrics. These models fall into three classes, which include FM-based methods (FM [29], NFM [30]), VAE-based methods (MultiVAE [31], MacridVAE [32], MacridVAE+FM), and causal-based methods (CausPref [12], COR [11]).
The comparison results are demonstrated in Table. II. Observations drawn from the results are as follows:
1) Our proposed DT3OR could achieve SOTA recommendation performance compared to classical FM-based recommendation algorithms. We infer that the reason behind this is that these methods do not design specific strategy to handle the user/item feature shift;
2) FM-based recommendation algorithms have a significant advantage on datasets with notable feature shifts, resulting in better performance compared to VAE-based algorithms on synthetic and Yelp datasets;
3) The test-time training with SSL tasks in our method leads to improved performance compared to causal-based recommendation methods.
In summary, our proposed DT3OR outperforms several recommendation methods across various datasets in the OOD scenario. Taking the results on meituan dataset for example, DT3OR exceeds the second-best method by margins of 12.77%, 0.87%, 11.88%, and 3.70% in R@50, R@100, N@50, and N@100, respectively. The test-time-training with dual SSL tasks could assist the model adapt to the shift user/item features in test-time phase under OOD environment, thus alleviating the influence of data shift.
V-D (RQ2) Parameter Sensitivity
@@
V-D1 Sensitivity Analysis for trade-off $\alpha$
@@
1) When $\alpha$ is assigned extreme values, such as $0.01$ or $100$, the performance of the recommendation system tends to decrease.
2) Extreme values can disrupt the balance between different loss components, whereas values around $1.0$ tend to yield better results.
@@
V-D2 Sensitivity Analysis for cluster number $K$
@@
1) The model achieves the best recommendation performance when $K$ is around 4. When $K$ takes extreme values, e.g., $K=100$, the performance will decrease dramatically. We speculate that this is because the interest centers become too scattered, making it difficult to accurately reflect the true preferences of users.
2) A similar situation occurs when $K=2$, where having too few interest centers fails to effectively capture the diverse preferences of users.
V-D3 Sensitivity Analysis for the threshold $\tau$
@@
V-E (RQ3) Ablation Studies
@@
1) We utilize clustering task to obtain the interest centers of different types of users, which could better benefit the model to learn the preferences (user-like-item) in test-time phase of shift user/item features.
@@
2) Benefiting from the data mining capacity, the contrastive mechanism could improve the discriminative capacity of the model, thus reducing the impact of the shift in user/item features.
V-F (RQ4) Visualization Analysis
@@
To reveal the intrinsic interest clustering structure, we utilize visualization experiments to display the distribution of the learned embeddings in this subsection. To be concrete, experiments are implemented with $t$-SNE algorithm [33] on the Synthetic dataset. We perform visualization experiments with varying cluster numbers, i.e., $K\in[3,4,5,6]$. As presented in Figure 9, the visual results emphasize that our proposed DT3OR method effectively identifies and represents the underlying interest clusters.
VI Related Work
VI-A Test-time Training
@@
Test-time training (TTT)[17, 34, 35] is a method to improve model generalization in recent years, which allows for partial adaptation of a model based on test samples. TTT addresses the issue of distribution shifts between the training and test sets by updating the model using self-supervised learning (SSL) tasks during the test-time phase. Initially, TTT was introduced to update the model by jointly optimizing the supervised loss and SSL loss during the test period. However, these paradigms necessitate the simultaneous optimization of the supervised loss and SSL loss. To surmount this constraint, Tent [36] presents a fully test-time training approach that solely uses test samples and a pre-trained model. Inspired by the success of TTT in handling out-of-distribution problems, we apply this technique to the OOD recommendation scenario. By leveraging TTT, we can adapt the recommendation model using test samples and a pre-trained model, improving its performance in the presence of distribution shifts.
VI-B Self-Supervised Learning
@@
Self-supervised learning (SSL) algorithms have recently garnered significant attention in the domain of graph neural networks (GNNs). SSL is a widely employed technique for training neural networks as evidenced in various studies [37, 38, 19, 39, 40]. A fundamental technique in SSL is contrastive learning, which focuses on increasing the similarity of positive pairs while decreasing the similarity of negative pairs. This approach has achieved remarkable success in graph learning [41, 42, 43, 44]. For instance, DCRN [45] strengthens feature decorrelation by aligning the representations of identical nodes across varied views while distancing the representations of dissimilar nodes in different views. Similarly, GMI [46] and HDMI [47] extend DGI [48] by incorporating edges and node features, to mitigate representation collapse. Additionally, the efficacy of amplifying mutual information to extract and learn embeddings for graph classification tasks has been demonstrated by MVGRL [49] and InfoGraph [50]. Building on these principles, GraphCL [51] and GRACE [52] produce two augmented views for SSL, and learn the representations by aligning the same nodes in both views and separating distinct nodes.
VI-C Recommender System
@@
Recommender systems [53, 54, 35] have become a vital tool for discovering user’s interests and preferences in various online platforms, which assume that the user and item follow the independent and identically distributed (I.I.D.). OOD recommendation has received limited attention, which can be roughly divided into three classes. 1) Disentangled recommendation approaches [15, 16] aim to learn factorized representations of user preferences to enhance robustness against distribution shifts. 2) Causal-based methods [11, 12] leverage causal learning techniques to address OOD problems. However, these methods often require interventions during the training process, which may be impractical when only a pre-trained model is available. 3) Model retraining [8] focuses on adapting the model to the OOD environment. This approach necessitates a trade-off between retraining frequency and computational cost. Benefiting from the success of test-time training (TTT), we introduce the TTT strategy to assist model adaptation in OOD scenarios in this paper.
VII Conclusion

### v2

@@ VI-A2 Evaluation Metrics
VI-B Experiment Settings
VI-B1 Implementation Details
The experiments are executed utilizing an NVIDIA 3090 GPU on the PyTorch platform. For the benchmark models, the performance values from the COR [19] paper are directly reported for all datasets. In our proposed approach, we utilized COR as the primary network and trained our network using the Adam optimizer [41]. The batch size was established at 500. Moreover, the training during test-time was carried out for 10 epochs. The trade-off hyper-parameter $\alpha$ was selected from the set $\{0.01,0.1,1.0,10,100\}$. The number of clusters $K$ was selected within the range of $\{2,4,10,20,100\}$, and the threshold $\tau$ for selecting high confidence samples was selected from $\{0.1,0.3,0.5,0.7,0.9\}$. The detailed hyper-parameter settings are provided in Tab. V.
VI-B2 Comparison Recommendation Methods
To substantiate the efficiency of our DT3OR, we contrast its performance with several baseline models, including FM [42], NFM [43], MultiVAE [44], MacridVAE [45], MacridVAE+FM, CausPref [20], COR [19], InvCF [46], PopGo [47], CausalDiffRec [48], DR-GNN [49]. Specifically,
@@ VI-B2 Comparison Recommendation Methods
  • DR-GNN [49] designed a distributionally robust optimization method to GNN-based recommendation algorithm.
  • PopGo [47] proposed a debiasing strategy to reduce the interaction-wise popularity shortcut. It learned a shortcut model to yield a shortcut degree of a user-item pair based on representations.
  • CausalDiffRec [48] addressed the out-of-distribution problem with causal diffusion method. By eliminating environmental confounding factors and learning invariant graph representations, the generalization ability of CausalDiffRec is improved for OOD data.
  • InvCF [46] designed an invariant collaborative filtering method to discover disentangled representations. It could reveal the latent preference and the popularity semantics.
Table IV: OOD recommendation performance comparison on two datasets with 11 baselines. Especially, the best results are emphasized in bold. Besides, $\dagger$ denotes results are statistically significant where the p-value is less than 0.05.
Dataset Amazon-Book Steam
Method R@50 R@100 N@50 N@100 R@50 R@100 N@50 N@100
FM ICDM 2010 2.19 3.62 0.47 0.98 9.52 12.08 2.47 3.57
NFM SIGIR 2017 2.64 3.89 0.51 1.03 9.04 11.85 2.26 2.68
MultiVAE WWW 2018 2.56 3.57 0.5 1.11 10.54 12.62 3.84 4.76
MacridVAE NeurIPS 2019 2.98 4.05 0.67 1.26 11.47 12.99 4.04 4.89
MacridVAE+FM NeurIPS 2019 3.04 4.86 1.01 1.28 12.75 13.27 5.04 5.87
CausPref WWW 2022 3.84 5.00 1.12 1.31 12.84 15.75 5.17 5.98
COR WWW 2022 4.12 6.57 1.24 2.35 13.44 22.64 5.72 7.13
InvCF WWW 2023 4.32 6.47 1.28 2.49 14.64 23.16 6.85 8.52
PopGo TOIS 2024 4.01 6.25 1.19 2.23 13.07 20.46 4.93 6.10
CausalDiffRec Preprint 2024 4.65 6.98 1.57 2.63 14.27 22.21 5.97 7.92
DR-GNN WWW 2024 4.23 6.17 1.35 2.48 13.15 21.24 5.06 6.08
Ours 5.07$\dagger$ 7.56$\dagger$ 2.26$\dagger$ 2.85$\dagger$ 15.72$\dagger$ 24.17$\dagger$ 7.65$\dagger$ 9.73$\dagger$
%Improvement 9.03% 8.31% 4.08% 8.37% 7.38% 4.36% 11.68% 14.20%
VI-C (RQ1) Performance Comparisons
We evaluate the experiments against 11 baseline models based on Recall and NDCG metrics. These models fall into four classes, which include FM-based methods (FM [42] and NFM [43]), VAE-based methods (MultiVAE [44], MacridVAE [45], and MacridVAE+FM), graph-based methods (DR-GNN [49]), causal learning methods (InvCF [46], CausPref [20], COR [19], CausalDiffRec [48], and PopGo [47]).
The comparison results are demonstrated in Tab. III. Observations drawn from the results are as follows:
  • Our proposed DT3OR could achieve SOTA recommendation performance compared to classical FM-based recommendation algorithms. We infer that the reason behind this is that these methods do not design specific strategy to handle the user/item feature shift;
  • FM-based recommendation algorithms have a significant advantage on datasets with notable feature shifts, resulting in better performance compared to VAE-based algorithms on synthetic and Yelp datasets;
  • Compared with graph-based and causal-based recommendation algorithms, our DT3OR demonstrates superior recommendation performance in out-of-distribution (OOD) scenarios. We attribute this improvement to the incorporation of test-time training with self-supervised learning (SSL) tasks, which enables our method to better adapt to shifted data distributions, leading to better performance.
In summary, our proposed DT3OR outperforms several recommendation methods across different scale datasets with different OOD scenario. Taking the results on meituan dataset for example, DT3OR exceeds the second-best method by margins of 12.77%, 0.87%, 11.88%, and 3.70% in R@50, R@100, N@50, and N@100, respectively. The test-time-training with dual SSL tasks could assist the model adapt to the shift user/item features in test-time phase under OOD environment, thus alleviating the influence of data shift.
VI-D (RQ2) Parameter Sensitivity
@@
VI-D1 Sensitivity Analysis for trade-off $\alpha$
@@
  • When $\alpha$ is assigned extreme values, such as $0.01$ or $100$, the performance of the recommendation system tends to decrease.
  • Extreme values can disrupt the balance between different loss components, whereas values around $1.0$ tend to yield better results.
@@
VI-D2 Sensitivity Analysis for cluster number $K$
@@
  • The model achieves the best recommendation performance when $K$ is around 4. When $K$ takes extreme values, e.g., $K=100$, the performance will decrease dramatically. We speculate that this is because the interest centers become too scattered, making it difficult to accurately reflect the true preferences of users.
  • A similar situation occurs when $K=2$, where having too few interest centers fails to effectively capture the diverse preferences of users.
VI-D3 Sensitivity Analysis for the threshold $\tau$
@@
Synthetic
Meituan
Figure 11: Sensitivity analysis of the hyper-parameter temperature $T$ on Synthetic and Meituan datasets.
VI-D4 Sensitivity Analysis for the temperature $T$
We conduct the experiments to verify the sensitivity of the temperature parameter in sharpen function, i.e., Eq. (5). To be specifically, we select the value of $T$ from the range of $\{0.1,0.3,0.5,0.7,0.9\}$. The results are presented in Fig. 11.
We could observe that lowering the temperature could achieve better recommendation performance in out-of-distribution scenario. We analyze the reason is that when $T\to 0$, the output of the Sharpen function will approximate to a one-hot distribution. The entropy will decrease with lowering temperature. The specific setting of $T$ is shown in Tab. V.
VI-E (RQ3) Ablation Studies
@@ VI-D4 Sensitivity Analysis for the temperature $T$
  • We utilize clustering task to obtain the interest centers of different types of users, which could better benefit the model to learn the preferences (user-like-item) in test-time phase of shift user/item features.
@@
  • Benefiting from the data mining capacity, the contrastive mechanism could improve the discriminative capacity of the model, thus reducing the impact of the shift in user/item features.
VI-F (RQ4) Visualization Analysis
@@
To reveal the intrinsic interest clustering structure, we utilize visualization experiments to display the distribution of the learned embeddings in this subsection. To be concrete, experiments are implemented with $t$-SNE algorithm [50] on the Synthetic dataset. We perform visualization experiments with varying cluster numbers, i.e., $K\in[3,4,5,6]$. As presented in Fig. 10, the visual results emphasize that our proposed DT3OR method effectively identifies and represents the underlying interest clusters.
VI-G (RQ5) Time and Space Cost
In this subsection, we conduct experiments to evaluate the time and space complexity using five datasets, i.e., Synthetic, Meituan, Yelp, Amazon, and Steam. The compared baselines are InvCF [46], COR [19], DR-GNN [49], and ours DT3OR. The experiments are conducted on the PyTorch deep learning platform with the 24G NVIDIA 3090. For the baselines, we utilized their source code available at Github. We measured time and space costs in seconds and gigabytes (GB), respectively. The results are reported by the average value over 10 runs. The observations are as follows.
  • Time Cost: The experimental results for time cost are presented in Fig.12. In this experiment, DT3OR utilizes COR[19] as the backbone network. We measured the training time for one epoch, averaged over 10 runs, with results reported in seconds. It can be observed that DT3OR’s training time is comparable to that of other algorithms.
  • Space Cost: The space cost experimental results are shown in Fig. 12. The results are reported as the average value over 10 runs and measured in gigabytes (GB). We observed that the memory costs of our DT3OR are acceptable when compared to other algorithms. Furthermore, these findings validate the effectiveness of the theoretical complexity analysis discussed in Sub-Section. IV-B.
Figure 12: Time cost and space cost on five datasets. Time and space complex is measured by seconds and gigabytes (GB), respectively. All results are the average value over 10 runs. The first row and second row correspond to time cost and space cost, respectively.
VI-H (RQ6) Stability Analysis
We implement experiments to explore the stability of DT3OR with different scales of distribution shift on Meituan and Amazon-Book datasets. We spit the dataset to construct temporal shift with timestamp, i.e., weekday as IID, weekend as OOD. By adjusting the proportion of the iterations with timestamp, specifically by modifying the number of weekend days, we could obtain different size of distribution shift. The experimental results are demonstrated in Fig. 13. We could observe as follows.
  • As the proportion of distribution shift increases, indicated by the number of weekend days, the performance of all methods deteriorates. This can be attributed to the models’ insufficient ability to handle large-scale distribution shifts, resulting in poor recommendation performance.
  • Our method consistently achieves better performance in most cases, highlighting the generalizability of the test-time training strategies.
Meituan
Amazon-Book
Figure 13: Different size of distribution shift on Meituan and Amazon-Book datasets. The x-axis represents the number of weekend days, while the y-axis denotes the R@50 performance.
VII Related Work
VII-A Test-time Training
@@
Test-time training (TTT) [51, 52] is a method to improve model generalization in recent years, which allows for partial adaptation of a model based on test samples. TTT addresses the issue of distribution shifts between the training and test sets by updating the model using self-supervised learning (SSL) tasks during the test-time phase. Initially, TTT was introduced to update the model by jointly optimizing the supervised loss and SSL loss during the test period. However, these paradigms necessitate the simultaneous optimization of the supervised loss and SSL loss. To surmount this constraint, Tent [53] presents a fully test-time training approach that solely uses test samples and a pre-trained model. Inspired by the success of TTT in handling out-of-distribution problems, we apply this technique to the OOD recommendation scenario. By leveraging TTT, we can adapt the recommendation model using test samples and a pre-trained model, improving its performance in the presence of distribution shifts.
VII-B Self-Supervised Learning
@@
Self-supervised learning (SSL) algorithms have recently garnered significant attention in the domain of graph neural networks (GNNs) [54, 55]. SSL is a widely employed technique for training neural networks as evidenced in various studies [56, 57, 58, 59, 60]. A fundamental technique in SSL is contrastive learning, which focuses on increasing the similarity of positive pairs while decreasing the similarity of negative pairs. This approach has achieved remarkable success in graph learning [61, 62]. For instance, DCRN [63] strengthens feature decorrelation by aligning the representations of identical nodes across varied views while distancing the representations of dissimilar nodes in different views. Similarly, GMI [64] extends DGI [65] by incorporating edges and node features, to mitigate representation collapse. Additionally, the efficacy of amplifying mutual information to extract and learn embeddings for graph classification tasks has been demonstrated by MVGRL [66] and InfoGraph [67]. Building on these principles, GraphCL [68] produces two augmented views for SSL, and learn the representations by aligning the same nodes in both views and separating distinct nodes.
VII-C Recommender System
@@
Recommender systems have become a vital tool for discovering user’s interests and preferences in various online platforms, which assume that the user and item follow the independent and identically distributed (I.I.D.). OOD recommendation has received limited attention, which can be roughly divided into three classes. 1) Disentangled recommendation approaches [25] aim to learn factorized representations of user preferences to enhance robustness against distribution shifts. 2) Causal-based methods [19, 20] leverage causal learning techniques to address OOD problems. However, these methods often require interventions during the training process, which may be impractical when only a pre-trained model is available. 3) Model retraining [27] focuses on adapting the model to the OOD environment. This approach necessitates a trade-off between retraining frequency and computational cost. Benefiting from the success of test-time training (TTT), we introduce the TTT strategy to assist model adaptation in OOD scenarios in this paper.
VIII Conclusion
@@
In this paper, we make an initial attempt to alleviate the out-of-distribution recommendation problem using a test-time training strategy. The key idea is to design an unsupervised self-distillation task and a contrastive learning task for the test-time model adaptation. The core operation in these tasks is obtaining the users’ preference centers through clustering. How to obtain high-quality clustering of preference centers when the dataset contains many user and item types is a valuable research question. Moreover, determining the optimal number of test-time training iterations for effective parameter updates remains a challenging issue. Excessive iterations may cause the model to lose valuable information learned during the initial training phase, while insufficient iterations may prevent the model from fully capturing the shifted data distribution. In the future, we will further explore the aforementioned issues to enhance the generalization ability of recommendation models in OOD scenarios.
Acknowledgments
This work was supported by the National Science and Technology Innovation 2030 Major Project under Grant No. 2022ZD0209103, the National Natural Science Foundation of China (project No. 62325604, 62276271, 62406336, 62376039, U24A20253, 62476281), and the Program of China Scholarship Council (No. 202406110009).
IX Appendix
IX-A Hyper-parameter Settings
Table V: Hyper-parameter setting for our method.
Dataset Synthetic data Meituan Yelp Amazon-Book Steam
Learning Rate 1e-4 1e-5 1e-4 1e-4 1e-4
Drop Rate 0.5 0.5 0.5 0.5 0.5
Epoch 10 10 10 10 10
$\alpha$ 1.0 1.0 1.0 1.0 1.0
$\tau$ 0.9 0.8 0.9 0.9 0.9
$T$ 0.1 0.3 0.3 0.1 0.1
$K$ 4 5 4 4 5
In this subsection, we report the statistics summary and hyper-parameter settings of our proposed method in Tab. V.
IX-B The Construction for Synthetic Data
We elaborate details on the process of constructing the Synthetic dataset in this part. Following the methodology in COR [19], we constructed the dataset through four primary steps:
1) Sampling user and item features;
2) Estimating user preferences;
3) Sampling user interactions;
4) Collecting OOD (Out-of-Distribution) data.
Specifically, the first step involves sampling user and item features from the standard Gaussian distributions, denoted as $\mathcal{N}(0,1)$ for users and $\mathcal{N}(-1,1)$ for items. We set the number of users and items to be 1,000 each. The second step pertains to user preference estimation, which is grounded in prior knowledge. For instance, higher income might positively influence preferences for more expensive items. The relationships can be either positive or negative. We calculate user preferences by summing the user features with either positive or negative weights and then apply a sigmoid function to introduce non-linearity into these relationships.
In the third step, sampling of user interactions is carried out. Utilizing the acquired user preferences and features of item, we ascertain the relevance of user-item interactions via the sigmoid function. Interaction data is subsequently sampled from the Bernoulli distribution [69]. The final step concerns the collection of OOD data. In this regard, we re-sample user characteristics from $\mathcal{N}(1,1)$ to accumulate user interactions in an OOD environment, maintaining item features constant. Steps 2 and 3 are then reiterated to sample new user interactions under these altered conditions. This rigorous process ensures a comprehensive and well-structured Synthetic dataset for our experiments.